# RISAR: RIS-assisted Human Activity Recognition with Commercial Wi-Fi Devices

### 27 Feb 2024 v1

@@ III-C Data Analysis and Preprocessing
Given that low-cost commercial Wi-Fi cards typically introduce diverse types of noise into CSI measurements [39], it necessitates a preprocessing phase for the CSI data. This preprocessing is vital to extract relevant features, substantially enhancing the effectiveness of algorithmic training and thus significantly contributing to the achievement of Wi-Fi sensing objectives.
The preprocessing phase of the RISAR system encompasses two critical components: the denoising and conversion of CSI data, followed by signal extraction. Initially, phase offset in CSI measurement values, including sampling time offset (STO) and sampling frequency offset (SFO), is addressed through a linear transformation method, as suggested in previous research [40]. This approach effectively mitigates shifts caused by STO/SFO, yielding more accurate phase values. Subsequently, our paper applies this method to ascertain precise phase values.
@@ III-C1 Activity feature analysis
Raw CSI measurements primarily capture amplitude and phase changes over time but lack inherent frequency components. The fast Fourier transform (FFT) is a standard technique for converting CSI measurements from the time to the frequency domain, aiding in power spectral density (PSD) extraction but compromising time-domain information [41, 42]. In contrast, the short-time Fourier transform (STFT) and discrete wavelet transform (DWT) enable the capture of features in both time and frequency domains [16, 43, 44, 17]. Fig. 3 illustrates the STFT (spectrogram) of CSI for various activities, revealing that activities with significant movements, such as walking, display high energy in the spectrogram. This crucial time-frequency information is instrumental in recognizing specific human activities.
@@
III-C2 Data denoising
CSI data is inherently replete with noise and often lacks the distinctness required to differentiate between various activities accurately. Consequently, the filtration of noise and extraction of salient features for classification through machine learning techniques is paramount. Prior research, including that by [45], has validated the effectiveness of PCA in this realm. PCA functions by reducing the dimensionality of extensive datasets, focusing primarily on key features that capture the bulk of signal information. Nevertheless, the efficacy of PCA is notably compromised in NLOS settings, where multipath fading and random noise are prevalent. Additionally, the CSI data exists in a high-dimensional vector space, characterized by its time series nature; it exhibits temporal variations ($T$ sampling packets) in conjunction with spatial variations ($N$ subcarriers). Extracting statistical information, particularly the temporal-spatial correlations, from these datasets presents a formidable challenge, one that conventional mathematical tools are ill-equipped to handle. To address this complexity, random matrix theory offers a mathematically rigorous solution for managing such data. In light of this, we introduce a novel denoising and feature extraction methodology: the high-dimensional factor model (HDFM), which is grounded in the Marčenko-Pastur (M-P) law [46].
The mathematical framework of HDFM is outlined as follows: Let $N$ represent the number of subcarriers, $T$ be the number of time sampling points, and $\mathbf{R}_{it}$ be the amplitude/phase of the $i$-th subcarrier at time $t$, where $i=1,2,\cdots,N$, $t=1,2,\cdots,T$, and $\frac{N}{T}\rightarrow c>0$. The CSI amplitude/phase matrix is expressed by Eq. (1), where $p$ is the number of temporal factors, $\mathbf{F}_{jt}$ denotes the $j$-th temporal factor at time $t$, and $\mathbf{L}_{ij}$ represents the loading of the $j$-th temporal factor on the $i$-th cross-sectional element. Furthermore, $\mathbf{U}_{it}$ stands for the additive random noise or idiosyncratic component.
@@ III-C2 Data denoising
where $\mathbf{L}^{(p)}\mathbf{F}^{(p)}$ represents the estimated common factor obtained from $p$ principal components. The HDFM method lies in examining the eigenvalue distribution of the covariance matrix of $p$-level residuals $\mathbf{U}_{it}^{(p)}$:
@@
Upon accurately extracting temporal factors, the residual spectral density $\mathbf{C}_{it}^{(p)}$ closely aligns with the M-P law, and these factors are utilized as input to the classification algorithm. The pseudocode for the proposed method is delineated in Algorithm 1. For a comprehensive understanding, the proof of the HDFM is elaborated in Appendix B.
@@ III-D Learning-based Algorithms
In our research, we integrate both machine learning (ML) and deep learning (DL) methodologies to evaluate the HDFM. We employ well-established ML algorithms like random forest (RF) [47], extreme gradient boosting (XGBoost) [48], and support vector machine (SVM) [49], alongside a novel dual-stream spatial-temporal attention deep learning framework designed for human activity classification. This framework’s architecture, as depicted in Fig. 4, begins with the preprocessing of CSI data through a spatial-temporal extractor. Following this, the extracted features are input into an attention layer, which utilizes an attention matrix to highlight the significance of various features and time steps. The integration of these features with the attention matrix elements results in a feature matrix with attention weights, which is then flattened into a feature vector. This vector facilitates the differentiation of action labels through a softmax classification layer.
@@
In the architecture proposed, the network incorporates three convolutional layers followed by a LSTM layer, which serves to further refine the feature extraction process. As delineated in Table II, detailed specifications for each layer are provided. Specifically, the convolutional layers are configured with 32, 64, and 128 neurons, respectively, while the LSTM layer is equipped with 512 hidden units. The variable $M$ denotes the batch size, and $K$ signifies the number of labels.
@@ IV Experiments and Results
where TP, TN, FP, and FN denote true positive, true negative, false positive, and false negative, respectively. True positive signifies the accurate identification of positive instances by the model while true negative represents correctly identified negative instances. False positive denotes the count of negative instances erroneously classified as positive, and false negative indicates the count of positive instances incorrectly classified as negative by the model.
@@ IV-A Performance Evaluation of RIS-enabled HAR
In this investigation, a comparative assessment is conducted to evaluate the efficacy of various algorithms within two distinct settings: environments without RIS configuration and those with RIS-optimized configurations, as elaborated in Table III. This examination highlights substantial enhancements in classifier accuracy following environmental optimization via RIS, with our proposed DS-STAN experiencing a significant accuracy uplift from 79.17% in non-configurable settings to 90.83% in RIS-optimized conditions. Comparable improvements are observed across other methodologies upon RIS activation, validating the critical role of RIS in augmenting the performance of activity-sensing technologies. Such advancements underscore the potential of RIS in revolutionizing wireless sensing systems, offering a pathway to achieve heightened precision and reliability in outcomes.
@@ IV-B Performance Evaluation of HDFM
In the preprocessing phase for CSI, PCA [17] and HDFM techniques are utilized for noise reduction and feature extraction. The extracted amplitude and phase features are subsequently input into various classifiers, including RF, XGBoost, SVM, and the proposed DS-STAN model that integrates time-frequency domain features. A comprehensive evaluation reveals HDFM’s significant enhancement of activity recognition accuracy in two datasets, with noticeable improvements over PCA-based approaches, as detailed in Table IV. Notably, the implementation of HDFM significantly enhances recognition accuracy in the Meeting room dataset, with improvements observed between 5.14% and 5.70% over PCA methods. Similarly, for the Office dataset, HDFM outperforms PCA, indicating accuracy gains ranging from 6.83% to 9.48%.
@@ IV-E Performance Evaluation with Other Existing Methods
In our comparative analysis, presented in Table V, the performance of our proposed DS-STAN method for activity recognition, utilizing the Meeting room dataset, significantly outperforms existing benchmark classifiers, achieving an exceptional accuracy rate of 96.39%. This evidence not only affirms the superior efficacy of our approach but also highlights its potential to emerge as a forefront solution in the HAR domain.
@@ V Conclusion
Our research presents RISAR, an innovative device-free passive human activity recognition system leveraging Wi-Fi signals, tailored for NLOS environments. RISAR significantly enhances the accuracy of activity recognition. The system’s architecture adeptly integrates an HDFM for effective signal extraction and noise reduction, alongside an advanced dual-stream spatial-temporal attention network model. The experimental results from diverse environments demonstrate the system’s proficiency. Notably, RISAR shows remarkable improvements in recognition accuracy, evident in the transition from PCA-based methods to HDFM techniques. The introduction of an attention layer within the deep learning framework further augments this accuracy, underscoring the importance of feature selection and emphasis in complex signal environments.
In conclusion, RISAR represents a significant advancement in human activity recognition technology, offering enhanced accuracy and adaptability in various settings. Its ability to operate effectively in challenging NLOS scenarios and its adaptability to different environmental conditions mark it as a promising solution for future IoT applications in smart homes, healthcare, and beyond. This study not only contributes to the field of wireless sensing and activity recognition but also paves the way for future research exploring the integration of RIS and advanced machine learning techniques in other domains.

### 20 Mar 2024 v3

@@ III-C Data Analysis and Preprocessing
The utilization of low-cost Wi-Fi cards for data collection introduces inherent imperfections in the CSI data, underscoring the necessity for preliminary data cleaning processes [36]. Moreover, the substantial dimensionality inherent to each data frame necessitates the development of data compression algorithms to enhance the efficiency of model training processes. In the context of learning-based recognition algorithms, the process of dimension reduction serves as a critical intermediary feature extraction technique.
@@ III-C1 Activity feature analysis
The principle that CSI encapsulates the cumulative effects of scattering, fading, and power decay within the channels is well-recognized. Consequently, the occurrence of varied human activities induces detectable fluctuations in the CSI, forming the basis for activity recognition methodologies. For enhanced elucidation of this phenomenon, the application of the short-time Fourier transform (STFT) on the CSI data is employed to delineate features across both temporal and frequency domains [6, 7]. Fig. 4 showcases the STFT (spectrogram) of the CSI corresponding to diverse activities, elucidating that actions characterized by pronounced movements, such as walking, manifest significantly elevated energy levels in the spectrogram. This pivotal time-frequency data proves essential for the accurate recognition of distinct human activities.
@@
III-C2 Data cleaning and compression
To address low-level imperfections such as phase offset in CSI measurements, which include sampling time offset (STO) and sampling frequency offset (SFO), we employ a linear transformation method [37]. This method effectively mitigates shifts caused by STO/SFO, resulting in more accurate phase values.
Within the domain of data compression, previous research [38] has validated the effectiveness of principal component analysis. One disadvantage of using PCA for compression lies in its assumption of linearity within the dataset. However, the CSI measurements often contain nonlinear patterns. Furthermore, the CSI data exhibit temporal correlations ($T$ sampling packets) in conjunction with spatial correlations ($N$ subcarriers). PCA may not adequately capture these complex dependencies within the data.
To address the temporal-spatial correlations in the CSI measurements, random matrix theory offers a mathematically rigorous solution. We propose a novel denoising and feature extraction algorithm known as the high-dimensional factor model (HDFM), which is based on the Marčenko-Pastur (M-P) law [39], one of the most remarkable laws in the random matrix theory.
Recall that $N$ represents the number of subcarriers and $T$ is the number of time sampling points. Let $\mathbf{R}_{it}$ denote the amplitude/phase of the $i$-th subcarrier at time $t$. Under the Kolmogorov condition, i.e., $T,N\to\infty$ and $\frac{N}{T}\to c>0$, the CSI amplitude/phase matrix can be represented by Eq. (1). Here $\mathbf{F}_{jt}$ denotes the $j$-th temporal factor at time $t$, $\mathbf{L}_{ij}$ represents the loading of the $j$-th temporal factor on the $i$-th cross-sectional element, $\mathbf{U}_{it}$ stands for the additive noise, and $p$ is the number of temporal factors.
@@ III-C2 Data cleaning and compression
where $\mathbf{L}^{(p)}\mathbf{F}^{(p)}$ represents the estimated common factor obtained from $p$ principal components.
In the proposed HDFM method, we examine the eigenvalue distribution of the covariance matrix of $p$-level residuals $\mathbf{U}_{it}^{(p)}$:
@@
The assumption is that if the factors are accurately eliminated, the residual spectral density $\mathbf{C}_{it}^{(p)}$ closely conforms to the Marčenko-Pastur law. This enables us to isolate the main factors present in the CSI data. For instance, by setting $p=8$, we compress the original CSI data frame from $270\times 5000$ to $8\times 5000$. The pseudocode for the proposed method is shown in Algorithm 1. For an in-depth proof of the HDFM, please refer to Appendix B.
@@ III-D Learning-based Algorithms
In this subsection, we integrate both machine learning (ML) and deep learning (DL) methodologies to evaluate the HDFM. We employ well-established ML algorithms like random forest (RF) [40], extreme gradient boosting (XGBoost) [41], and support vector machine (SVM) [42], alongside a novel dual-stream spatial-temporal attention deep learning framework designed for human activity classification.
This framework’s architecture, as depicted in Fig. 5, begins with the preprocessing of CSI data through a spatial-temporal extractor. Following this, the extracted features are input into an attention layer, which utilizes an attention matrix to highlight the significance of various features and time steps. The integration of these features with the attention matrix elements results in a feature matrix with attention weights, which is then flattened into a feature vector. This vector facilitates the differentiation of action labels through a softmax classification layer.
@@
In the architecture proposed, the network incorporates three convolutional layers followed by an LSTM layer, which serves to further refine the feature extraction process. As delineated in Table II, detailed specifications for each layer are provided. Specifically, the convolutional layers are configured with 32, 64, and 128 neurons, respectively, while the LSTM layer is equipped with 512 hidden units. The variable $M$ denotes the batch size, and $K$ signifies the number of labels.
@@ IV Experiments and Results
where TP, TN, FP, and FN denote true positive, true negative, false positive, and false negative, respectively.
@@ IV-A Performance Evaluation of RIS-enabled HAR
In this subsection, a comparative assessment is conducted to evaluate the efficacy of various algorithms within two distinct settings: environments without RIS configuration and those with RIS-optimized configurations, as elaborated in Table III. This examination highlights substantial enhancements in classifier accuracy following environmental optimization via RIS, with our proposed DS-STAN experiencing a significant accuracy uplift from 79.17% in non-configurable settings to 90.83% in RIS-optimized conditions.
Comparable improvements are observed across other methodologies upon RIS activation, validating the critical role of RIS in augmenting the performance of activity-sensing technologies. Such advancements underscore the potential of RIS in revolutionizing wireless sensing systems, offering a pathway to achieve heightened precision and reliability in outcomes.
@@ IV-B Performance Evaluation of HDFM
In the preprocessing phase for CSI, PCA [7] and HDFM techniques are utilized for noise reduction and feature extraction. The extracted amplitude and phase features are subsequently input into various classifiers, including RF, XGBoost, SVM, and the proposed DS-STAN model that integrates time-frequency domain features. A comprehensive evaluation reveals HDFM’s significant enhancement of activity recognition accuracy in two datasets, with noticeable improvements over PCA-based approaches, as detailed in Table IV. Notably, the implementation of HDFM significantly enhances recognition accuracy in the Meeting room dataset, with improvements observed between 5.14% and 5.70% over PCA methods. Similarly, for the Office dataset, HDFM outperforms PCA, indicating accuracy gains ranging from 6.83% to 9.48%. It validates that the features calculated through our proposed HDFM can be more representative of the CSI data than those from PCA, which demonstrates the superiority of HDFM.
@@ IV-E Performance Evaluation with Other Existing Methods
In our comparative analysis, presented in Table V, the performance of our proposed DS-STAN method for activity recognition, utilizing the Meeting room dataset, significantly outperforms existing methods including LSTM, ABLSTM, Wi-multi, DeepSeg, LCED, CeHAR, and CNN-GRU-CNN, achieving an exceptional accuracy rate of 96.39%. The result not only affirms the superior efficacy of our approach but also highlights its potential to emerge as a forefront solution in the HAR domain.
@@ V Conclusion
This paper presents RISAR, a novel passive RIS-assisted human activity recognition system leveraging Wi-Fi signals, particularly optimized for low signal-to-noise ratio environments. By integrating a high-dimensional factor model for signal extraction and noise reduction with a dual-stream spatial-temporal attention network, RISAR significantly outperforms traditional PCA-based methods in accuracy and efficiency, achieving up to 97.26% accuracy in experimental settings. This system’s robustness, demonstrated across diverse environmental conditions, underlines its potential for wide-ranging IoT applications, from smart homes to healthcare. The study not only contributes substantially to the fields of wireless sensing and activity recognition but also paves the way for future exploration into combining RIS technology with advanced machine learning techniques, marking a significant step forward in realizing more intuitive and seamless smart environments.